could someone please help me with this problem?
(3-i) / (1+i) I know you use the conjugate and multiply by (3-i) / (1+I) I think,but I don't think I did something right.Thanks for your help in advance.
Sincerely,LH

Found 2 solutions by ptaylor, stanbon:
Answer by ptaylor(2198)


OK, YOU ARE ABSOLUTELY RIGHT: SO LETS USE THE CONJUGATE AND WHAT THIS WILL DO IS GET RID OF THE RADICAL IN THE DENOMINATOR (AS YOU KNOW, THE CONJUGATE OF 1+i IS 1-i.
((3-i)(1-i))/(1+i)(1-i))
The numerator reduces to(using FOIL: First, Outer, Inner, Last):
3-3i-i+i^2 (i^2=-1) so we have for the numerator:
3-4i-1=2-4i
For the denominator we have:
1-i+i-i^2; (-(i^2))=(-(-1))=+1 so we have
1+1=2
Putting the numerator and denominator back together, we have:
(2-4i)/2 divide numerator and denominator by 2 we get:
1-2i-------------------ans

Hope this helps----ptaylor

Answer by stanbon(75887)

(3-i) / (1+i)
---------------
Multiply numerator and denominator by (1-i) to get:
(3-i)(1-i)/2
= [3-1-4i]/2
= [-4-4i]/2
= [-2-2i]
===============
Cheers,
Stan H.